# Language-Specific Neurons: The Key to Multilingual Capabilities in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) demonstrate remarkable multilingual capabilities without being pre-trained on specially curated multilingual parallel corpora. It remains a challenging problem to explain the underlying mechanisms by which LLMs process multilingual texts. In this paper, we delve into the composition of Transformer architectures in LLMs to pinpoint language-specific regions. Specially, we propose a novel detection method, language activation probability entropy (LAPE), to identify language-specific neurons within LLMs. Based on LAPE, we conduct comprehensive experiments on two representative LLMs, namely LLaMA-2 and BLOOM. Our findings indicate that LLMs’ proficiency in processing a particular language is predominantly due to a small subset of neurons, primarily situated in the models’ top and bottom layers. Furthermore, we showcase the feasibility to “steer” the output language of LLMs by selectively activating or deactivating language-specific neurons. Our research provides important evidence to the understanding and exploration of the multilingual capabilities of LLMs.
@@
Tianyi Tang1,3 * †thanks: * This work was done during internship at MSRA., Wenyang Luo1 †thanks: † Equal contribution., Haoyang Huang2, Dongdong Zhang2 Xiaolei Wang1, Wayne Xin Zhao1,3 🖂thanks: 🖂 Corresponding author , Furu Wei2, Ji-Rong Wen1,3 1 Gaoling School of Artificial Intelligence, Renmin University of China 2 Microsoft Research Asia, China 3 Beijing Key Laboratory of Big Data Management and Analysis Methods steventianyitang@outlook.com  luowenyang@ruc.edu.cn  wxl1999@foxmail.com {haohua,dozhang,fuwei}@microsoft.com    batmanfly@gmail.com
@@ 2.2 Language Activation Probability Entropy
Our research primarily focuses on pre-trained base models (e.g., LLaMA-2 and BLOOM), rather than fine-tuned models that have undergone instruction tuning or RLHF (e.g., LLaMA-2-Chat and BLOOMZ), which helps reduce other influencing factors. Specially, we feed existing LLMs with multilingual texts, each written in a single language. For the $j$-th neuron in the $i$-th layer, we then compute the activation probability when processing texts in language $k$:
$p^{k}_{i,j}=\mathbb{E}\left(\mathbb{I}(\text{act\_fn}(\hat{\bm{h}}^{i}\bm{W}^{i}_{1})_{j}>0)\mid\text{language }k\right),$ (4)
@@ 3.1 Experimental Setup
We conduct the study based on two public LLMs: LLaMA-2 Touvron et al. (2023b), primarily pre-trained on English texts and known for its excellence as a language foundational model, and BLOOM Scao et al. (2022), known for its multilingual proficiency with a more balanced language distribution. We investigate multiple versions of LLaMA-2, specifically the 7B, 13B, and 70B models, containing approximately 352K, 553K, and 2.29M neurons, respectively. For BLOOM, we select the 7.1B version, consisting of roughly 492K neurons. The languages selected include English (en), Simplified Chinese (zh), French (fr), Spanish (es), Vietnamese (vi), Indonesian (id), and Japanese (ja). We exclude ja for BLOOM since it has not been pre-trained on Japanese copora.
@@
$\bullet$ Open-ended generation: To evaluate model’s multilingual generation capabilities in real-world scenarios, we translate a set of questions from the Vicuna dataset Chiang et al. (2023) into the target languages. These questions span a broad spectrum of topics, excluding mathematics and coding queries to maintain a focus on language processing proficiency, avoiding confounding variables. We utilize greedy search with repetition penalty of 1.1 to generate output and then employ GPT-4 to assess the output text on a scale ranging from 1 to 10 following Zheng et al. (2023).
@@
(b) Language activation value entropy (LAVE): This is a variant of our proposed method, wherein we substitute the activation probability with the mean activation value across languages. Similarly, we normalize the activation value and calculate the entropy to find neurons with high activation value in response to particular languages.
(c) Parameter variation (PV): By extending the work of Zhao et al. (2023a), we compare the model parameters before and after monolingual instruction tuning to identify language-specific parameters. In particular, we train individual models on the Alpaca instruction datasets Taori et al. (2023) and its multilingual version Chen et al. (2023b) which comprise 52,000 instances for each target. These models undergo training for two epochs, with a batch size of 128 and a constant learning rate of 1e-5. We mainly consider the parameters inside the MHA and FFN modules, i.e., the weight matrices in Equations 12, and 3. We compute the rate of change across various languages, and select parameters that exhibit a low rate of change in one or two languages but a high rate in others. In detail, we refine the change ratio by subtracting the maximum value and then conduct L1 normalization for entropy calculation.
(d) Random selection (RS): Additionally, we add a baseline to randomly select neurons for each language, serving as a reference for different methods.
@@ 3.2 Main Perturbation Experiments
(b) LLaMA-2 (13B)
(c) LLaMA-2 (70B)
(d) BLOOM (7B)
@@
Figure 2 presents the perturbation results (measured by PPL change on language modeling task) of different identification methods on LLaMA-2. Overall, we can see that large impact values for LAPE mainly occur in diagonal entries. It indicates that our LAPE method is adept at identifying language-specific neurons. Deactivating neurons associated with a specific language predominantly affects the PPL results of that language, with negligible effects on others. In contrast, the variant utilizing activation values (LAVE) causes cross-lingual interference (e.g., the entry $\langle id,es\rangle$), and the rest two methods PV and RS fails to exhibit clear language-specific patterns.
@@
Table 1: Performance of LLaMA-2 (70B) on the multilingual Vicuna as evaluated by GPT-4. The “normal” row is baseline scores without deactivation. Subsequent rows are scores with deactivation of specific neurons.
We further investigate whether our method is effective across different model sizes and model types. As depicted in Figure 3, the language-specific degradation patterns remain evident for three LLaMA models (7B, 13B, 70B) and a BLOOM model (7B). An interesting find is that neurons in larger models tend to be specialized for a single language rather than being shared among two or more languages. Furthermore, we can find that there exists a high correlation between Chinese and Japanese: when we deactivating the neurons specific to one language, the performance of the other language would be affected (e.g., the entries $\langle zh,ja\rangle$ and $\langle ja,zh\rangle$). By inspecting into the identified neurons of the two languages, we note that a substantial amount of neurons (approximately 25%) actually overlap for both languages. It is likely because Chinese and Japanese partially share common characters. In addition, it can be observed that our LAPE method leads to similar findings on BLOOM-7B (Figure 3(d)). These results further verify the effectiveness of generality of our proposed LAPE method.
@@
我是一個登上珠穆朗瑪峰頂峰的登山者。 I am a mountaineer who has climbed to the top of Mount Everest. 當我站在珠my朗ma峰頂峰,我感到非常興奮和欣慰。 …
@@ 3.3.2 Structural Distribution Analysis
In Figure 5, we present the layer-wise distribution of language-specific neurons across various layers, which has a skewed “U” shape. This finding indicates that language-specific neurons have a pronounced concentration in the top and bottom layers. Specifically, the second layer is composed of approximately 7,000 neurons, while layers 5 through 47 only contain about 100 neurons each. Further, the neuron count gradually increases, with the final four layers each comprising over 1,000 neurons. The complete statistics about the layer-wise distribution across various languages are reported in Table 10 of the Appendix.
@@
To understand why such a skewed distribution occurs, we seek explanations from multilingual semantic representation by exploring how multilingual aligned texts are represented across the layers. Specially, we employ the multilingual Vicuna dataset (Section 3.1), comprising of aligned texts in different languages. Given a group of aligned texts, we feed them into the LLM and obtain the sentence embeddings for each layer. We then compute the mean sentence embedding similarity (SES) among the aligned texts across language in Figure 6. Interestingly, the SES curve shows an opposite trend with the distribution of language-specific neurons. At the beginning, the similarity quickly increases, then reaches the peak, and gradually decreases to a small value. This finding shows: at the bottom layers, the LLM needs to map aligned texts of different languages into the shared representation space, thus requiring more language-specific neurons for semantic transformation; while at top layers serving for token generation, the LLM needs to handle vocabulary mapping, which also requires the support of more language-specific neurons.
@@ 3.3.3 Language Dominance Analysis
To compute the dominance degree, we still use the multilingual Vicuna dataset (Section 3.1). From the results of LLaMA-2 (70B) in Figure 6(a), we can see that the mean SES is obviously higher than all other languages when the target language is English. As English is the high-resource language of LLaMA-2, low-resource languages need to be aligned with English for achieving better performance. When it comes to BLOOM (170B) in Figure 6(b), several languages (e.g., English and Chinese) shows dominance, since it is originally pre-trained on multilingual corpora.
@@ 3.3.4 Case Study
Researchers have observed that when prompting in one language, language models may generate responses in a different language, such a phenomenon referred to as the off-target language issue Gu et al. (2019); Sennrich et al. (2023). We speculate that language-specific neuron might not be activated in this case. Thus, we attempt to manually activating these neurons by increasing their activation value to the average activation value of this language. As shown in Table 4, it is evident that the model demonstrates a significantly higher propensity to respond in the expected language, as indicated by the language accuracy. Additionally, the generation quality, as assessed by GPT-4 metrics, shows remarkable improvement.
@@
?‘Cómo puedo mejorar mis habilidades de gestión del tiempo?
@@ 4 Related Work
Neuron analysis has gained much attention in recent years following neurobiological human brain Friederici (2011); Parr et al. (2022). Originating from vision models Bau et al. (2020); Mu and Andreas (2020), neuron analysis views the neuron activation as the recall of learned knowledge Sajjad et al. (2022). Researchers widely adopt them to analyze the source of some specific abilities or skills language models, including sentiment analysis (Radford et al., 2017), knowledge storing (Dai et al., 2022), task solving Wang et al. (2022). Further, some recent studies have also found that certain neurons can convey special contexts Gurnee et al. (2023); Bills et al. (2023), such as positions Voita et al. (2023) and linguistic related properties Bau et al. (2018); Xin et al. (2019); Dalvi et al. (2019, 2020). Moreover, Gurnee et al. (2024) use the Pearson correlation to calculate neuron similarity and find some universal neurons across models. In contrast, neuron analysis is not widely utilized in multilingual areas, and we extend this method to LLMs and unveil the mechanism of their multilingual ability.

### v2

@@ Abstract
Large language models (LLMs) demonstrate remarkable multilingual capabilities without being pre-trained on specially curated multilingual parallel corpora. It remains a challenging problem to explain the underlying mechanisms by which LLMs process multilingual texts. In this paper, we delve into the composition of Transformer architectures in LLMs to pinpoint language-specific regions. Specially, we propose a novel detection method, language activation probability entropy (LAPE), to identify language-specific neurons within LLMs. Based on LAPE, we conduct comprehensive experiments on several representative LLMs, such as LLaMA-2, BLOOM, and Mistral. Our findings indicate that LLMs’ proficiency in processing a particular language is predominantly due to a small subset of neurons, primarily situated in the models’ top and bottom layers. Furthermore, we showcase the feasibility to “steer” the output language of LLMs by selectively activating or deactivating language-specific neurons. Our research provides important evidence to the understanding and exploration of the multilingual capabilities of LLMs.
@@
Tianyi Tang1 * †thanks: * This work was done during internship at MSRA., Wenyang Luo1 †thanks: † Equal contribution., Haoyang Huang2, Dongdong Zhang2 Xiaolei Wang1, Wayne Xin Zhao1 🖂thanks: 🖂 Corresponding author , Furu Wei2, Ji-Rong Wen1,3 1 Gaoling School of Artificial Intelligence, Renmin University of China 2 Microsoft Research Asia, China 3 School of Information, Renmin University of China {steventianyitang,wengyang_luo}@outlook.com   wxl1999@foxmail.com {haohua,dozhang,fuwei}@microsoft.com   batmanfly@gmail.com
@@ 2.2 Language Activation Probability Entropy
Our research primarily focuses on pre-trained foundation models (e.g., LLaMA-2 and BLOOM), rather than fine-tuned models that have undergone instruction tuning or RLHF, which helps reduce other influencing factors. Specially, we feed existing LLMs with multilingual texts, each written in a single language. For the $j$-th neuron in the $i$-th layer, we then compute the activation probability when processing texts in language $k$:
$p^{k}_{i,j}=\mathbb{E}\left(\mathbb{I}(\text{act\_fn}(\tilde{\bm{h}}^{i}\bm{W}^{i}_{1})_{j}>0)\mid\text{language }k\right),$ (4)
@@ 3.1 Experimental Setup
We conducted our study primarily on two publicly available large language models (LLMs): LLaMA-2 Touvron et al. (2023b) and BLOOM Scao et al. (2022). Among them, LLaMA-2 is recognized for its excellence as a foundational model, primarily pre-trained on English texts, while BLOOM is noted for its multilingual proficiency due to a balanced distribution of training languages. Specifically, we investigate multiple versions of LLaMA-2: the 7B, 13B, and 70B models, which contain approximately 352K, 553K, and 2.29M neurons, respectively. For BLOOM, we select the 7.1B version, consisting of roughly 492K neurons. The languages we focus on include English (en), Simplified Chinese (zh), French (fr), Spanish (es), Vietnamese (vi), Indonesian (id), and Japanese (ja). We exclude Japanese (ja) for BLOOM since it has not been pre-trained on Japanese corpora. To verify the generality of our method, we also include LLMs under different settings, including LLaMA-2 Chat, OPT Zhang et al. (2022), Mistral Jiang et al. (2023), and Phi-2 Javaheripi et al. (2023).
@@
$\bullet$ Open-ended generation: To evaluate the model’s multilingual generation capabilities in real-world scenarios, we translate a set of questions from the Vicuna dataset Chiang et al. (2023) into target languages using gpt-4-0125-preview. The questions span a broad spectrum of topics, deliberately excluding mathematics and coding queries to maintain a focus on language processing proficiency and avoid confounding variables. We utilize greedy search with a repetition penalty of 1.1 to generate output. The resulting texts are then assessed by GPT-4 on a scale ranging from 1 to 10, following the methodology described by Zheng et al. (2023).
@@
(b) Language activation probability (LAP): There are also methods to identify neurons directly. But most of them are infeasible due to the high computational cost Gurnee et al. (2024); Dai et al. (2022). Inspired by Voita et al. (2023), we apply their method by identifying a neuron as language-specific if its activation probability exceeds 95%.
(c) Language activation value entropy (LAVE): This is a variant of our proposed method, wherein we substitute the activation probability with the mean activation value across languages. Similarly, we normalize the activation value and calculate the entropy to find neurons with high activation value in response to particular languages.
(d) Parameter variation (PV): By extending the work of Zhao et al. (2023a), we compare the model parameters before and after monolingual instruction tuning to identify language-specific parameters. In particular, we train individual models on the Alpaca instruction datasets Taori et al. (2023) and its multilingual version Chen et al. (2023b) which comprise 52,000 instances for each target. These models undergo training for two epochs, with a batch size of 128 and a constant learning rate of 1e-5. We mainly consider the parameters inside the MHA and FFN modules, i.e., the weight matrices in Equations 12, and 3. We compute the rate of change across various languages, and select parameters that exhibit a low rate of change in one or two languages but a high rate in others. In detail, we refine the change ratio by subtracting the maximum value and then conduct L1 normalization for entropy calculation.
(e) Random selection (RS): Additionally, we add a baseline to randomly select neurons for each language, serving as a reference for different methods as shown in Figure 8 in Appendix.
@@ 3.2 Main Perturbation Experiments
(b) LLaMA-2-Chat (7B)
(c) LLaMA-2 (13B)
(d) LLaMA-2 (70B)
(e) BLOOM (7B)
(f) OPT (6.7B)
(g) Mistral (7B)
(h) Phi-2 (2.7B)
@@
Figure 2 presents the perturbation results (measured by PPL change on language modeling task) of different identification methods on LLaMA-2. Overall, we can see that large impact values for LAPE mainly occur in diagonal entries. It indicates that our LAPE method is adept at identifying language-specific neurons. Deactivating neurons associated with a specific language predominantly affects the PPL results of that language, with negligible effects on others. In contrast, the variant utilizing activation values (LAVE) causes cross-lingual interference (e.g., the entry $\langle id,es\rangle$), and the rest methods fail to exhibit clear language-specific patterns.
@@
Table 1: Performance of LLaMA-2 (70B) on the multilingual Vicuna as evaluated by GPT-4. The “normal” row is baseline scores without deactivation while the “random” row is with random deactivation. Subsequent rows are scores with deactivation of specific neurons.
We further investigate whether our method is effective across different model sizes and model types. As depicted in Figure 3, the language-specific degradation patterns are evident for LLaMA-2, BLOOM, OPT, Mistral, and Phi-2. An interesting find is that neurons in larger models tend to be specialized for a single language rather than being shared among two or more languages. Furthermore, we can find that there exists a high correlation between Chinese and Japanese: when we deactivating the neurons specific to one language, the performance of the other language would be affected (e.g., the entries $\langle zh,ja\rangle$ and $\langle ja,zh\rangle$). By inspecting into the identified neurons of the two languages, we note that a substantial amount of neurons (approximately 25%) actually overlap for both languages. It is likely because Chinese and Japanese partially share common characters. In addition, it can be observed that our LAPE method leads to similar findings on BLOOM and other LLMs, which further verify the generality of our proposed LAPE method.
@@
{CJK*}UTF8bsmi我是一個登上珠穆朗瑪峰頂峰的登山者。 I am a mountaineer who has climbed to the top of Mount Everest. 當我站在珠my朗ma峰頂峰,我感到非常興奮和欣慰。 …
@@ 3.3.2 Structural Distribution Analysis
In Figure 5, we present the layer-wise distribution of language-specific neurons across various layers, which has a skewed “U” shape. This finding indicates that language-specific neurons have a pronounced concentration in the top and bottom layers. Specifically, the second layer has approximately 7,000 language-specific neurons, while layers 5 through 47 only contain about 100 neurons each. Further, the neuron count gradually increases, with the final four layers each comprising over 1,000 neurons. The complete statistics about the layer-wise distribution across various languages are reported in Table 10.
@@
To understand why such a skewed distribution occurs, we seek explanations from multilingual semantic representation by exploring how multilingual aligned texts are represented across the layers. Specially, we employ the multilingual Vicuna dataset (Section 3.1), comprising of aligned texts in different languages. Given a group of aligned texts, we feed them into the LLM and obtain the sentence embedding of each text for each layer. We then compute the mean sentence embedding similarity (SES) between each pair of the aligned texts across languages in Figure 6. Interestingly, the SES curve shows an opposite trend with the distribution of language-specific neurons. At the beginning, the similarity quickly increases, then reaches the peak, and gradually decreases to a small value. This finding suggests that: at the bottom layers, the LLM needs to map aligned texts of different languages into the shared representation space, thus requiring more language-specific neurons for semantic transformation; while at top layers serving for token generation, the LLM needs to handle vocabulary mapping, which also requires the support of more language-specific neurons.
@@ 3.3.3 Language Dominance Analysis
To compute the dominance degree, we still use the multilingual Vicuna dataset (Section 3.1). From the results of LLaMA-2 (70B) in Figure 7(a), we can see that the mean SES is obviously higher than all other languages when the target language is English. As English is the high-resource language of LLaMA-2, low-resource languages need to be aligned with English for achieving better performance. When it comes to BLOOM (170B) in Figure 7(b), several languages (e.g., English and Chinese) show dominance, since it is originally pre-trained on multilingual corpora.
@@ 3.3.4 Case Study
Researchers have observed that when prompting in one language, language models may generate responses in a different language, such a phenomenon referred to as the off-target language issue Gu et al. (2019); Sennrich et al. (2023). We speculate that language-specific neuron might not be activated in this case. Thus, we attempt to manually activate these neurons by increasing their activation value to the average activation value of this language. As shown in Table 4, it is evident that the model demonstrates a significantly higher propensity to respond in the expected language, as indicated by the language accuracy. Additionally, the generation quality, as assessed by GPT-4 metrics, shows remarkable improvement.
@@
¿Cómo puedo mejorar mis habilidades de gestión del tiempo?
@@ 4 Related Work
Neuron analysis has gained significant attention in recent years, paralleling research in neurobiological studies of the human brain Friederici (2011); Parr et al. (2022). Originating from vision models Bau et al. (2020); Mu and Andreas (2020), neuron analysis views neuron activation as the recall of learned knowledge Sajjad et al. (2022). Researchers widely adopt these methods to analyze the sources of specific abilities or skills in language models, including sentiment analysis (Radford et al., 2017), knowledge storage (Dai et al., 2022), and task-solving Wang et al. (2022).
Recent studies have also discovered that certain neurons can convey specialized contexts Gurnee et al. (2023); Bills et al. (2023), such as positional information Voita et al. (2023) and linguistic properties Bau et al. (2018); Xin et al. (2019); Dalvi et al. (2019, 2020). Moreover, Gurnee et al. (2024) utilize Pearson correlation to calculate neuron similarity, identifying some universal neurons across models. In contrast to previous research, we have developed a method applicable to LLMs that unveils the mechanism of their multilingual abilities. This approach offers a more practical and effective solution for neuron analysis in multilingual scenarios.
@@ 5 Conclusion
Acknowledgement
This work was partially supported by National Natural Science Foundation of China under Grant No. 62222215, Beijing Natural Science Foundation under Grant No. L233008 and 4222027. Xin Zhao is the corresponding author.